Circular No. 6355

Central Bureau for Astronomical Telegrams INTERNATIONAL ASTRONOMICAL UNION

Postal Address: Central Bureau for Astronomical Telegrams Smithsonian Astrophysical Observatory, Cambridge, MA 02138, U.S.A. IAUSUBS@CFA.HARVARD.EDU or FAX 617-495-7231 (subscriptions)

BMARSDEN@CFA.HARVARD.EDU or DGREEN@CFA.HARVARD.EDU (science) Phone 617-495-7244/7440/7444 (for emergency use only)

COMET C/1996 B2 (HYAKUTAKE) M. Festou, Observatoire Midi-Pyrenées, Toulouse; P. Feldman, Johns

Hopkins University, for the IUE ToO Team for Comet Observations; and R. Gonzalez and P. M. Rodriguez, European Space Agency IUE Observatory, report: “Observations of the comet were made on Mar. 23.12 and 24.15 UT with the International Ultraviolet Explorer (IUE) using the new One-Gyro mode of operation, which had to be implemented after a major spacecraft problem on Mar. 6.084. The operational modifications associated with this limit the duration of exposures on moving targets. Low-dispersion spectra were recorded with both cameras, both with the nucleus in and out of the aperture. The LWP spectrum (range 200–330 nm) reveals the presence of strong OH emission, from which a water-production rate of 3.5 × 10^{29} s^{−1} is inferred (r = 1.05 AU). This indicates a sevenfold increase from the water-production rate measured by Schleicher et al. (IAUC 6344), while the heliocentric distance decreased by only ∼ 0.10 AU. Emission features usually seen in LWP range, such as those due to C_{2} (231.3 nm), the CS bands, and the CO_{2}^{+} B-X and NH 336.0-nm bands, also appear on top of a strong continuum. The SWP spectrum (115–199 nm) shows the presence of weak carbon and oxygen emissions. The fourth positive band system of CO is identified by its bands longward of the 0-0 band, the strongest being 0-1 at 159.7 nm. The intensity distribution of the bands indicates a significant opacity of the CO atmosphere. The optical thickness of the 180.0-nm component of the S I 181.3-nm triplet is very large. It is likely that many cometary emissions seen throughout the spectrum will be optically thick (OH, CN, etc.).”

Naked-eye m_{1} and tail-length estimates: Mar. 24.10 UT, +0.1, 30^{o} (K. Černis, Moletai, Lithuania); 24.33, −0.5, 70^{o} (C. S. Morris, Lockwood Valley, CA); 25.58, +0.9, ∼ 80^{o} (S. J. O’Meara, Volcano, HI); 26.00, −0.4, 90^{o} (A. Baransky, Kiev, Ukraine); 26.34, +0.3, ∼ 100^{o} (J. V. Scotti, Corona de Tucson, AZ); 26.58, +0.9, 100^{o} (O’Meara).

COMET C/1996 B1 (SZCZEPANSKI) Total visual magnitude and coma-diameter estimates: Mar. 8.81 UT,

7.2, 18′ (K. Hornoch, Lelekovice, Czech Republic, 10×80 binoculars); 13.08, 8.0, 12′ (J. E. Bortle, Stormville, NY, 10×50 binoculars); 22.77, 9.3, 8′ (A. Baransky, Kiev, Ukraine, 0.11-m reflector); 25.18, 8.8, 8′ (C. E. Spratt, Victoria, BC, 0.10-m refractor).

1996 March 26  \begin{matrix}c \\ ◯ \end{matrix} Copyright 1996 CBAT  Daniel W. E. Green